An object of mass m travelling with speed v has a head on collision with another object of mass m travelling with speed v in the opposite direction. The two objects stick together after the collision.
What is the total loss of kinetic energy in the collision?
A)0
B)1/2mv^2
C)mv^2
D)2mv^2
1
Expert's answer
2020-11-16T07:41:05-0500

Conservation of momentum:


"mv-mv=2mu=0\\u=\frac{m}{s}"

"\Delta K=0.5mv^2+0.5mv^2-0=mv^2"

Answer: C) mv^2.